Consider U=q2+q2

Derive the Mashallian dd function from q1 and q2.


1
Expert's answer
2021-01-25T13:49:51-0500
"\frac {\delta u}{\delta q_1}=2q_1""\frac {\delta u}{\delta q_2}=2q_2"

"\frac {MU_1}{p_1}=\frac {MU_2}{p_2}"


"\frac {2q_1}{p_1}=\frac {2q_2}{p_2}"

"I=p_1q_1+p_2q_2"


"q_2=\frac {p_2}{p_1}q_1"


"q_1^*=\frac {Ip_1}{p_1^2+p_2^2}"


"q_2^*=\frac {Ip_2}{p_1^2+p_2^2}"